OpenStudy (anonymous):

Medal Given ∑[(k-1),5,(k-8)^3] Expand The Series

OpenStudy (anonymous):

@Loser66 HELPPPPPP

OpenStudy (loser66):

don't understand the notation, use draw box below to write it,please

OpenStudy (anonymous):

\[\sum_{k-1}^{5}\rightarrow(k-8)^3\]

OpenStudy (anonymous):

Expand it.

OpenStudy (loser66):

what does \(\rightarrow \) mean?

OpenStudy (anonymous):

i couldnt figure out how to put that on the side of sigma. but its just supposed to be on the side

OpenStudy (loser66):

ok, so far, we have 5 terms, 1,2,3,4,5 k =1, (k-8) =?

OpenStudy (anonymous):

-7

OpenStudy (loser66):

(-7)^3=?

OpenStudy (anonymous):

343

OpenStudy (loser66):

-343 not 343, ok?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and then what?

OpenStudy (loser66):

so, \(k_1=-343\) now, k =2; (k-8)=?

OpenStudy (anonymous):

-6\

OpenStudy (anonymous):

-6

OpenStudy (loser66):

(-6)^3=?

OpenStudy (anonymous):

-216?

OpenStudy (loser66):

yep, \(k_2 = -216\) so, keep up \(k_3,k_4,k_5\) then add them together, that's it

OpenStudy (anonymous):

i think i got it thanks!

OpenStudy (loser66):

hehe,... glad to hear. Didn't think that how easy it is, right?